OpenStudy (anonymous):

Ann and Billy both entered a quiz. The quiz had twenty questions and points were allocated as follows: • p points were added for each correctly answered question. • q points were deducted for each incorrect (or unanswered) question. Ann got 15 questions correct and scored 65 points. Billy got 11 questions correct and scored 37 points. Use matrices to find the value of q?

OpenStudy (gorv):

ann

OpenStudy (gorv):

correct = 15 points =15*p=15p wrong =5 points = 5*q so 15p-5q=65

OpenStudy (gorv):

billy correct=11 points=11p wrong=9 points=9q total 11p-9q=37

OpenStudy (gorv):

15p-5q=65 11p-9q=37 you have two equation now ..use matrix method and solve its simple now